# Wilsonian Renormalization of Neural Network Gaussian Processes

### v2

@@ IV Renormalization beyond the Gaussian
$\int\!\mathcal{D}\varphi_{>}$ $P[\varphi_{>}|\varphi_{<}]\,e^{-\frac{2\Phi_{<}^{\top}\Phi_{>}+\Phi_{>}^{\top}\Phi_{>}}{2\sigma^{2}}}$ (46)
$\approx\int\!\mathcal{D}\varphi_{>}P[\varphi_{>}|\varphi_{<}]\,e^{-\frac{\Phi_{<}^{\top}\Phi_{>}}{\sigma^{2}}}.$
@@
$\int\!\mathcal{D}\varphi_{>}$ $P[\varphi_{>}|\varphi_{<}]\,e^{-\frac{\Phi_{<}^{\top}\Phi_{>}}{\sigma^{2}}}$ (47)
@@
$\int$ $\!\mathcal{D}\varphi_{q}P[\varphi_{q}|\varphi_{<}]e^{-\frac{\Phi_{<}^{\top}\Phi_{q}}{\sigma^{2}}}$ (48)
$=\frac{1}{\sqrt{2\pi}}\int\!\mathcal{D}\varphi_{q}e^{-\frac{\varphi^{2}_{q}}{2}-\frac{\Phi_{<}^{\top}\Phi_{q}}{\sigma^{2}}}\left[1+A\varphi_{q}+B\left(\varphi_{q}^{2}-1\right)\right]$ (49)
$=\frac{1}{\sqrt{2\pi}}\int\!\mathcal{D}\varphi_{q}e^{-\frac{\varphi^{2}_{q}}{2}-\frac{1}{\sigma^{2}}\sum_{m}\Phi_{m<}\Phi_{mq}}$
×$\qquad\qquad\qquad\qquad\times\left[1+A\varphi_{q}+B\left(\varphi_{q}^{2}-1\right)\right]$ (50)
$\approx\exp\left[\frac{1+2B}{2\sigma^{4}}\left(\Phi_{<}^{\top}\bm{f}_{q}\right)^{2}-\frac{A}{\sigma^{2}}\Phi_{<}^{\top}\bm{f}_{q}\right]\leavevmode\nobreak\ ,$ (51)
@@
$S_{\mathrm{int}}=$ (53)
$-\eta\int\!\mathcal{D}\varphi P[\varphi]\,\exp\left[-\frac{1}{2\sigma^{2}}\left(\Phi_{<}^{\top}\Phi_{<}+2\Phi_{<}^{\top}\Phi_{>}+\Phi_{>}^{\top}\Phi_{>}\right)\right]$
$\approx-\eta\int\!\mathcal{D}\varphi_{<}\,P[\varphi_{<}]\,$
×$\quad\times\exp\left[-\frac{1}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}+\frac{1+2B}{2\sigma^{4}}\left(\Phi_{<}^{\top}\bm{f}_{q}\right)^{2}-\frac{A}{\sigma^{2}}\,\Phi_{<}^{\top}\bm{f}_{q}\right]\leavevmode\nobreak\ .$
@@
$Z^{M}\right>_{\eta}$ $=e^{-\eta}\int\!\mathcal{D}\bm{f}_{<}\,e^{-S_{0}[\bm{f}_{<}]}\int\!\mathcal{D}\bm{f}_{q}\,e^{-S_{0}[\bm{f}_{q}]-S_{\mathrm{int}}}\leavevmode\nobreak\ .\quad$ (54)
@@
$\int\!\mathcal{D}\bm{f}_{q}\,e^{-S_{0}[\bm{f}_{q}]-S_{\mathrm{int}}}$
$\approx\exp\left\{\eta\int\!\mathcal{D}\varphi_{<}\,P[\varphi_{<}]\,e^{-\frac{\Phi_{<}^{\top}\Phi_{<}}{2\sigma^{2}}+\lambda_{q}(1+2B)\frac{\Phi_{<}^{\top}\Phi_{<}}{2\sigma^{4}}}\right\}\leavevmode\nobreak\ ,\quad$ (55)
@@ IV.1 Functional RG flow
$\delta W_{>}(x)$ $=2\sum_{k_{1}\neq q^{\prime}}W_{k_{1}q^{\prime}}He_{2}[\varphi_{k_{1}},\varphi_{q^{\prime}}]$ (63)
$\qquad\qquad\qquad+W_{q^{\prime},q^{\prime}}He_{2}[\varphi_{q^{\prime}},\varphi_{q^{\prime}}]\leavevmode\nobreak\ .$
@@
$S$ int
$=$ $\int\!\mathcal{D}\varphi\,P[\varphi_{<},\varphi_{>}]\,$
×$\times\exp\!\left\{-\frac{1+\delta W_{<}+\delta W_{>}}{2\sigma^{2}}\left(\Phi_{<}^{\top}\Phi_{<}+\Phi_{>}^{\top}\Phi_{>}+2\Phi_{<}^{\top}\Phi_{>}\right)\right\}\leavevmode\nobreak$
$=$ $\int\!\mathcal{D}\varphi_{<}\,P[\varphi_{<}]\exp\!\left\{-\frac{1+\delta W_{<}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}\right\}\int\!\mathcal{D}\varphi_{>}P[\varphi_{>}|\varphi_{<}]$ (64)
×$\times\exp\!\left\{-\frac{1+\delta W_{<}+\delta W_{>}}{2\sigma^{2}}\left(\Phi_{>}^{\top}\Phi_{>}+2\Phi_{<}^{\top}\Phi_{>}\right)\right.$
$\qquad\qquad\qquad\qquad\left.-\frac{1}{2\sigma^{2}}\delta W_{>}\Phi_{<}^{\top}\Phi_{<}\right\}\leavevmode\nobreak\ ,$
@@
×$\times\int\!\mathcal{D}\varphi_{>}P[\varphi_{>}|\varphi_{<}]\exp\!\left\{-\frac{1+\delta W_{<}}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}\right\}\leavevmode\nobreak$
×$\times\left(1-\frac{\delta W_{>}}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}-\frac{\delta W_{>}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}\right)\leavevmode\nobreak\ .$
@@
$-\int\!\mathcal{D}\varphi_{>}{\cal N}[0,1;\varphi_{>}]\left(1-\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}+\frac{1}{2}\left[\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}\right]^{2}\right)$
×$\qquad\times\left(\frac{\delta W_{>}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}+\frac{\delta W_{>}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}\right),$ (66)
where in the first bracket we kept up to second order in the greater GP mode since, as seen before, higher terms would contribute as $\delta c$ to a power larger than $1$ following the greater GP mode integration. In the second bracket, we kept leading order in $U$.
The above integral leads to five terms, whose contributions to the renormalized GP action are analyzed in appendix C. The crucial observation that each such contribution appears as $O(\delta W_{>})O(\delta c/\sigma^{2})$ and consequently, as $\delta W_{>}$ is itself $O(\delta c/\sigma^{2})$— of order $(\delta c/\sigma^{2})^{2}$. Such contributions are negligible within our momentum-shell (i.e., leading order in $\delta c/\sigma^{2}$) treatment. Indeed $O(\delta c/\sigma^{2})^{2}$ terms, even after being summed $c/\delta c$-times during the RG process (where $c$ is the total variance of the integrated out modes), would be $O(c/\delta c)O(\delta c^{2})=O(\delta c)$ and vanish in the $\delta c\rightarrow 0$ limit. In contrast, the $O(\delta c)$ terms (“Wilsonian” terms) may contribute $O(c/\delta c)O(\delta c)=O(c)$ and hence remain finite as $\delta c\rightarrow 0$.
@@
$=\frac{1-\frac{2\,\delta c}{\sigma^{2}_{0}}B_{0}(x)}{\sigma_{\delta c}^{2}}-\frac{\delta c(1+2B_{\delta c}(x))}{\sigma^{4}_{\delta c}}+O(\delta c^{2})\leavevmode\nobreak\ .$ (67)
@@
which again integrates to 1 under $\mathrm{d}\mu_{x}$ (i.e., normalization is preserved).
This time, no new structure appeared in our RG step, hence we may iterate it $M$ times to obtain the following RG flow
$\sigma^{2}_{c}$ $=\sigma^{2}_{0}+c$ (69)
$W_{c}(x)$ $=1-\sum_{m=0}^{M-1}\delta cB_{m\delta c}(x)/\sigma_{c}$ (70)
[…]
where $\delta cM=c$ and the integral is obtain in the limit $\delta c\rightarrow 0$.
IV.2 A solvable toy example: single mode integration
@@
From a spectral bias perspective [5], the average predictor for this model is zero, since the target is orthogonal to all non-zero eigenspaces under the input measure. [^11] By this we mean that the inner product on the associated Hilbert space vanishes, $\int\!\mathrm{d}\mu_{x}\,K(x,y)He_{5}(x)=0$. This is not necessarily the case however, since the features ($\varphi$, defined in (20)) will be non-Gaussian unless $\phi(x)$ is an affine function of $x\sim\mathcal{N}[0,1]$. Statistically therefore, the more common events in which $x$ lies closer to zero may bias the predictor on small datasets to act as if it were seeing the function $-10x^{3}+15x$, which does have non-vanishing overlap with $He_{1}(x)$. We show below that our RG approach can accurately predict such effects, and that they can be viewed as stemming from an MSE loss weighted by $\left(1-\lambda_{2}\sigma^{-2}He_{2}(x)^{2}\right)$. Indeed, while $y=He_{5}(x)$ and $He_{1}(x)$ are not coupled by standard MSE loss, the extra weighting term yields a non-zero overlap via integrals of the type $\int\!\mathrm{d}\mu_{x}He_{1}(x)\left(\lambda_{2}\sigma^{-2}He_{2}(x)^{2}\right)He_{5}(x)$.
We shall first derive the above weighted form of the MSE loss, and then discuss its effect on predictions. Since the kernel mode associated with $He_{5}(x)$ is frozen/non-fluctuating, we do not treat it as a GP mode, and instead integrate out only the mode associated with $\lambda_{2}$. [^12] That is, we view $\bm{y}_{<}(x)$ as $\bm{y}_{1}He_{1}(x)+\bm{y}_{5}He_{5}(x)=He_{5}(x)$, with $y_{1}=0$ and $y_{5}=1$, and similarly $f_{<}(x)=f_{1}x$ with $f_{5}=0$. Alternatively, one may treat $y$ as part of the greater target modes, where its effect would then enter as a subleading shift to the lesser target. We denote this by $\varphi_{>}\eqqcolon\varphi_{2}$, and the associated replica-space GP mode by $\bm{f}_{2}$, with $\bm{y}_{2}=0$ (i.e., $q=2$ in our notation above). Furthermore, we take ${\bm{f}}_{1}$ to be the replica space vector associated with the linear (lesser) mode and (i.e., $He_{1}(x)$) and ${\bm{y}}_{1}$ to be the corresponding target. Converting to feature variables $\varphi_{<}\eqqcolon\varphi_{1}$ and $\varphi_{>}\eqqcolon\varphi_{2}$ via (22), we have
$P[\varphi_{1}]$ $=\int\mathrm{d}\mu_{x}\delta\left(\varphi_{1}-x\right)=\mathcal{N}[0,\mathbb{1};\varphi_{1}]\leavevmode\nobreak\ ,$ (74)
$P[\varphi_{1},\varphi_{2}]$ $=P[\varphi_{1}]P[\varphi_{2}|\varphi_{1}]$ (75)
@@
Next, we must integrate the greater GP mode $f_{2}$ in the replicated partition function (52), $Z^{M}\right>_{\eta}=e^{-\eta}\int\!\mathcal{D}\bm{f}\,e^{-S_{0}[\bm{f}]-S_{\mathrm{int}}}$. As in (55), we capitalize on the fact that $f_{2}$ scales as $\sqrt{\lambda_{2}}$, and keep all contributions below order $f_{2}^{4}=O(\lambda_{2}^{2})$. After expanding both exponentials, performing the integral, and re-exponentiating, we again find that the effect is to replace $f_{2}$ by the variance under $S_{0}[\bm{f}_{2}]$, [^13] Here we have also dropped the term $M\frac{\lambda_{q}}{2\sigma^{2}}He_{2}(\varphi_{1})^{2}$ that arises from $|\bm{f}_{q}|^{2}=\mathrm{tr}(\bm{f}_{q}^{\top}\bm{f}_{q})\rightarrow M\lambda_{q}$, since this vanishes in the replica limit. Such a factor of $M$ did not arise in (55), since there $\left(\Phi_{<}^{\top}\bm{f}_{q}\right)^{2}\sim\bm{f}_{q}\bm{f}_{q}^{\top}$ does not involve a trace over replica indices.
$Z^{M}\right>_{\eta}$ $=\int\!\mathcal{D}f_{1}\exp\bigg{\{}-S_{0}[\bm{f}_{1}]$ (77)
$\qquad+\eta\int\!\mathrm{d}\varphi_{1}P[\varphi_{1}]e^{-\frac{1}{2\sigma^{2}}\Phi_{1}^{\top}\Phi_{1}\left(1-\frac{\lambda_{2}}{\sigma^{2}}He_{2}(\varphi_{1})^{2}\right)}\bigg{\}}$
$=\int\!\mathcal{D}f_{1}\exp\bigg{\{}-S_{0}[\bm{f}_{1}]$
$\qquad+\eta\int\!\mathrm{d}\mu_{x}\,e^{-\frac{1}{2\sigma^{2}}\Phi_{1}^{\top}\Phi_{1}\left(1-\frac{\lambda_{2}}{\sigma^{2}}He_{2}(x)^{2}\right)}\bigg{\}}\leavevmode\nobreak\ ,$
where on the second line we used (75) to replace $\varphi_{1}$, $\int\!\mathrm{d}\varphi_{1}P[\varphi_{1}]$ with $x$, $\int\!\mathrm{d}\mu_{x}$, respectively. We thus observe the advertised reweighting of the MSE term, namely
@@
Next, we perform an EK-type approximation on (77). Namely, we assume $\sigma^{2}$ is large enough compared to $\left(\bm{f}_{1}(x)-\bm{y}_{1}(x)\right)$ to perform a Taylor expansion of the exponent appearing in the action. We note in passing that the next-to-leading order in this expansion [32] happens to vanish for $He_{n\geq 5}(x)$ (but not, e.g., for $He_{3}(x)$). Within the leading order expansion, the action in (77) becomes
@@
To test this empirically, we sampled various draws of datasets with $n=100$ from $x\sim{\cal N}[0,1]$, performed GP regression on these with the above kernel using a fixed random test set with $n_{\mathrm{test}}=1000$, and averaged the results over the various draws (see fig. 1). Doing so for $\lambda_{2}=0$ shows an almost zero outcome, as expected. However, taking finite $\lambda_{2}$ reveals a trend that spectral bias [32, 5] misses. Notably, the fact that our toy example involves only one mode which is integrated out, together with the need to take its variance to be small ($\lambda_{2}/\sigma^{2}\ll 1$) as well as make it unlearnable ($\eta\lambda_{2}/\sigma^{2}\ll 1$) results in a rather small effect. Still, it serves to demonstrate the conceptual phenomena whose importance depends on the setting or the desired accuracy.
Figure 1: Non-Gaussian features and spatial re-weighting effects. Theory versus experiment for the model of sec. IV.2 with $n=100$ datapoints $\sigma^{2}=400$ and $\lambda_{>}:=\lambda_{2}=0.1$ (unless stated otherwise). Learning a 5th Hermite polynomial, using a kernel capable of expressing only 1st and 2nd Hermite polynomials should give a zero average predictor (green line) based on the standard theory [5, 32]. However, due to spatial re-weighting, a coupling between 1st and 5th Hermite polynomial arises leading to a non-zero result.
V Discussion
In this work, we applied the powerful formalism of Wilsonian RG to Gaussian Process regression. In particular, treating the output of a trained DNN as a statistical ensemble, we split the associated fluctuations into high- and low-energy modes (based on their stiffness in a random network). Provided the variance of the higher modes is much smaller than the ridge parameter, these may be integrated out in a tractable manner leading to a renormalized GP regression problem. In simple settings where the features associated with those modes are Gaussian, this results in a renormalized ridge parameter. In more general settings, one obtains an effective GP regression problem where the ridge parameter becomes spatially dependent. We then applied this to a simple toy model, in which the effects determined via our RG approach are demonstrably missed in the standard eigen-learning framework [5].

### v3

@@ V Renormalization beyond the Gaussian
$\int\!\mathcal{D}\varphi_{>}P[\varphi_{>}|\varphi_{<}]\,e^{-\frac{2\Phi_{<}^{\top}\Phi_{>}+\Phi_{>}^{\top}\Phi_{>}}{2\sigma^{2}}}\approx\int\!\mathcal{D}\varphi_{>}P[\varphi_{>}|\varphi_{<}]\,e^{-\frac{\Phi_{<}^{\top}\Phi_{>}}{\sigma^{2}}}.$ (47)
@@
$\int\!\mathcal{D}\varphi_{q}$ $P[\varphi_{q}|\varphi_{<}]e^{-\frac{\Phi_{<}^{\top}\Phi_{q}}{\sigma^{2}}}=\frac{1}{\sqrt{2\pi}}\int\!\mathcal{D}\varphi_{q}e^{-\frac{\varphi^{2}_{q}}{2}-\frac{\Phi_{<}^{\top}\Phi_{q}}{\sigma^{2}}}\left[1+A\varphi_{q}+B\left(\varphi_{q}^{2}-1\right)\right]$ (49)
$=\frac{1}{\sqrt{2\pi}}\int\!\mathcal{D}\varphi_{q}e^{-\frac{\varphi^{2}_{q}}{2}-\frac{1}{\sigma^{2}}\sum_{m}\Phi_{m<}\Phi_{mq}}\left[1+A\varphi_{q}+B\left(\varphi_{q}^{2}-1\right)\right]$
$\approx\exp\left[\frac{1+2B}{2\sigma^{4}}\left(\Phi_{<}^{\top}\bm{f}_{q}\right)^{2}-\frac{A}{\sigma^{2}}\Phi_{<}^{\top}\bm{f}_{q}\right]\leavevmode\nobreak\ ,$
@@
$S_{\mathrm{int}}=$ $-\eta\int\!\mathcal{D}\varphi P[\varphi]\,\exp\left[-\frac{1}{2\sigma^{2}}\left(\Phi_{<}^{\top}\Phi_{<}+2\Phi_{<}^{\top}\Phi_{>}+\Phi_{>}^{\top}\Phi_{>}\right)\right]$ (51)
$\approx-\eta\int\!\mathcal{D}\varphi_{<}\,P[\varphi_{<}]\,\exp\left[-\frac{1}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}+\frac{1+2B}{2\sigma^{4}}\left(\Phi_{<}^{\top}\bm{f}_{q}\right)^{2}-\frac{A}{\sigma^{2}}\,\Phi_{<}^{\top}\bm{f}_{q}\right]\leavevmode\nobreak\ .$
@@
$Z^{M}\right>_{\eta}=e^{-\eta}\int\!\mathcal{D}\bm{f}_{<}\,e^{-S_{0}[\bm{f}_{<}]}\int\!\mathcal{D}\bm{f}_{q}\,e^{-S_{0}[\bm{f}_{q}]-S_{\mathrm{int}}}\leavevmode\nobreak\ .$ (52)
@@
$\int\!\mathcal{D}\bm{f}_{q}\,e^{-S_{0}[\bm{f}_{q}]-S_{\mathrm{int}}}\approx\exp\left\{\eta\int\!\mathcal{D}\varphi_{<}\,P[\varphi_{<}]\,e^{-\frac{\Phi_{<}^{\top}\Phi_{<}}{2\sigma^{2}}+\lambda_{q}(1+2B)\frac{\Phi_{<}^{\top}\Phi_{<}}{2\sigma^{4}}}\right\}\leavevmode\nobreak\ ,\quad$ (53)
@@ V.1 Functional RG flow
$\delta W_{>}(x)$ $=2\sum_{k_{1}\neq q^{\prime}}W_{k_{1}q^{\prime}}He_{2}[\varphi_{k_{1}},\varphi_{q^{\prime}}]+W_{q^{\prime},q^{\prime}}He_{2}[\varphi_{q^{\prime}},\varphi_{q^{\prime}}]\leavevmode\nobreak\ .$
@@
$S_{\mathrm{int}}=$ $\int\!\mathcal{D}\varphi\,P[\varphi_{<},\varphi_{>}]\,\exp\!\left\{-\frac{1+\delta W_{<}+\delta W_{>}}{2\sigma^{2}}\left(\Phi_{<}^{\top}\Phi_{<}+\Phi_{>}^{\top}\Phi_{>}+2\Phi_{<}^{\top}\Phi_{>}\right)\right\}\leavevmode\nobreak$
$=$ $\int\!\mathcal{D}\varphi_{<}\,P[\varphi_{<}]\exp\!\left\{-\frac{1+\delta W_{<}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}\right\}\int\!\mathcal{D}\varphi_{>}P[\varphi_{>}|\varphi_{<}]$
×$\times\exp\!\left\{-\frac{1+\delta W_{<}+\delta W_{>}}{2\sigma^{2}}\left(\Phi_{>}^{\top}\Phi_{>}+2\Phi_{<}^{\top}\Phi_{>}\right)-\frac{1}{2\sigma^{2}}\delta W_{>}\Phi_{<}^{\top}\Phi_{<}\right\}\leavevmode\nobreak\ ,$
@@
×$\times\int\!\mathcal{D}\varphi_{>}P[\varphi_{>}|\varphi_{<}]\exp\!\left\{-\frac{1+\delta W_{<}}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}\right\}\left(1-\frac{\delta W_{>}}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}-\frac{\delta W_{>}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}\right)\leavevmode\nobreak\ .$
@@
$-\int\!\mathcal{D}\varphi_{>}{\cal N}[0,1;\varphi_{>}]\left(1-\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}+\frac{1}{2}\left[\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}\right]^{2}\right)\left(\frac{\delta W_{>}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}+\frac{\delta W_{>}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}\right),$ (62)
where in the first bracket we kept up to second order in the greater GP mode since, as seen before, higher terms would contribute as $\delta c$ to a power larger than $1$ following the greater GP mode integration. In the second bracket, we kept up to leading order in $U$.
The above integral leads to five terms, whose contributions to the renormalized GP action are analyzed in appendix D. The crucial observation that each such contribution appears as $O(\delta W_{>})O(\delta c/\sigma^{2})$ and consequently, as $\delta W_{>}$ is itself $O(\delta c/\sigma^{2})$— of order $(\delta c/\sigma^{2})^{2}$. Such contributions are negligible within our momentum-shell (i.e., leading order in $\delta c/\sigma^{2}$) treatment. Indeed $O(\delta c/\sigma^{2})^{2}$ terms, even after being summed $c/\delta c$-times during the RG process (where $c$ is the total variance of the integrated out modes), would be $O(c/\delta c)O(\delta c^{2})=O(\delta c)$ and vanish in the $\delta c\rightarrow 0$ limit. In contrast, the $O(\delta c)$ terms (“Wilsonian” terms) may contribute $O(c/\delta c)O(\delta c)=O(c)$ and hence remain finite as $\delta c\rightarrow 0$.
@@
$=$ $\,\frac{1-\frac{2\,\delta c}{\sigma^{2}_{0}}B_{0}(x)}{\sigma_{\delta c}^{2}}-\frac{\delta c(1+2B_{\delta c}(x))}{\sigma^{4}_{\delta c}}+O(\delta c^{2})\leavevmode\nobreak\ .$
@@
which again integrates to 1 under $\mathrm{d}\mu_{x}$ (i.e., normalization is preserved). This time, no new structure appeared in our RG step, hence we may iterate it $M$ times to obtain the following RG flow
$\sigma^{2}_{c}$ $=\sigma^{2}_{0}+c\leavevmode\nobreak\ ,$ (64)
$W_{c}(x)$ $=1-\sum_{m=0}^{M-1}\delta cB_{m\delta c}(x)/\sigma_{c}$
$\approx 1-2\int_{0}^{c}\!\mathrm{d}c^{\prime}\,B_{c^{\prime}}(x)/\sigma_{c}^{2}\leavevmode\nobreak\ ,$
where $\delta cM=c$ and the integral is obtain in the limit $\delta c\rightarrow 0$. Note that the full, input-dependent ridge parameter is $W_{c}(x)/\sigma_{c}^{2}$, cf. (55); here, we have split this into the data independent component and the spatial reweighting.
V.2 A solvable toy example: single mode integration
@@
From a spectral bias perspective canatar , the average predictor for this model is zero, since the target is orthogonal to all non-zero eigenspaces under the input measure [^16] By this we mean that the inner product on the associated Hilbert space vanishes, $\int\!\mathrm{d}\mu_{x}\,K(x,y)He_{5}(x)=0$.. This is not necessarily the case however, since the features ($\varphi$, defined in (20)) will be non-Gaussian unless $\phi(x)$ is an affine function of $x\sim\mathcal{N}[0,1]$. Statistically therefore, the more common events in which $x$ lies closer to zero may bias the predictor on small datasets to act as if it were seeing the function $-10x^{3}+15x$, which does have non-vanishing overlap with $He_{1}(x)$. We show below that our RG approach can accurately predict such effects, and that they can be viewed as stemming from an MSE loss weighted by $\left(1-\lambda_{2}\sigma^{-2}He_{2}(x)^{2}\right)$. Indeed, while $y=He_{5}(x)$ and $He_{1}(x)$ are not coupled by standard MSE loss, the extra weighting term yields a non-zero overlap via integrals of the type $\int\!\mathrm{d}\mu_{x}He_{1}(x)\left(\lambda_{2}\sigma^{-2}He_{2}(x)^{2}\right)He_{5}(x)$.
We shall first derive the above weighted form of the MSE loss, and then discuss its effect on predictions. Since the kernel mode associated with $He_{5}(x)$ is frozen/non-fluctuating, we do not treat it as a GP mode, and instead integrate out only the mode associated with $\lambda_{2}$ [^17] That is, we view $\bm{y}_{<}(x)$ as $\bm{y}_{1}He_{1}(x)+\bm{y}_{5}He_{5}(x)=He_{5}(x)$, with $y_{1}=0$ and $y_{5}=1$, and similarly $f_{<}(x)=f_{1}x$ with $f_{5}=0$. Alternatively, one may treat $y$ as part of the greater target modes, where its effect would then enter as a subleading shift to the lesser target.. We denote this by $\varphi_{>}\eqqcolon\varphi_{2}$, and the associated replica-space GP mode by $\bm{f}_{2}$, with $\bm{y}_{2}=0$ (i.e., $q=2$ in our notation above). Furthermore, we take ${\bm{f}}_{1}$ to be the replica space vector associated with the linear (lesser) mode and (i.e., $He_{1}(x)$) and ${\bm{y}}_{1}$ to be the corresponding target. Converting to feature variables $\varphi_{<}\eqqcolon\varphi_{1}$ and $\varphi_{>}\eqqcolon\varphi_{2}$ via (21), we have
$P[\varphi_{1}]$ $=\int\mathrm{d}\mu_{x}\delta\left(\varphi_{1}-x\right)=\mathcal{N}[0,\mathbb{1};\varphi_{1}]\leavevmode\nobreak\ ,$
$P[\varphi_{1},\varphi_{2}]$ $=P[\varphi_{1}]P[\varphi_{2}|\varphi_{1}]$
@@
Next, we must integrate the greater GP mode $f_{2}$ in the replicated partition function (50), $Z^{M}\right>_{\eta}=e^{-\eta}\int\!\mathcal{D}\bm{f}\,e^{-S_{0}[\bm{f}]-S_{\mathrm{int}}}$. As in (53), we capitalize on the fact that $f_{2}$ scales as $\sqrt{\lambda_{2}}$, and keep all contributions below order $f_{2}^{4}=O(\lambda_{2}^{2})$. After expanding both exponentials, performing the integral, and re-exponentiating, we again find that the effect is to replace $f_{2}$ by the variance under $S_{0}[\bm{f}_{2}]$ [^18] Here we have also dropped the term $M\frac{\lambda_{q}}{2\sigma^{2}}He_{2}(\varphi_{1})^{2}$ that arises from $|\bm{f}_{q}|^{2}=\mathrm{tr}(\bm{f}_{q}^{\top}\bm{f}_{q})\rightarrow M\lambda_{q}$, since this vanishes in the replica limit. Such a factor of $M$ did not arise in (53), since there $\left(\Phi_{<}^{\top}\bm{f}_{q}\right)^{2}\sim\bm{f}_{q}\bm{f}_{q}^{\top}$ does not involve a trace over replica indices.,
$Z^{M}\right>_{\eta}$ $=\int\!\mathcal{D}f_{1}\exp\bigg{\{}-S_{0}[\bm{f}_{1}]+\eta\int\!\mathrm{d}\varphi_{1}P[\varphi_{1}]e^{-\frac{1}{2\sigma^{2}}\Phi_{1}^{\top}\Phi_{1}\left(1-\frac{\lambda_{2}}{\sigma^{2}}He_{2}(\varphi_{1})^{2}\right)}\bigg{\}}$ (69)
$=\int\!\mathcal{D}f_{1}\exp\bigg{\{}-S_{0}[\bm{f}_{1}]+\eta\int\!\mathrm{d}\mu_{x}\,e^{-\frac{1}{2\sigma^{2}}\Phi_{1}^{\top}\Phi_{1}\left(1-\frac{\lambda_{2}}{\sigma^{2}}He_{2}(x)^{2}\right)}\bigg{\}}\leavevmode\nobreak\ ,$
where on the second line we used (V.2) to replace $\varphi_{1}$, $\int\!\mathrm{d}\varphi_{1}P[\varphi_{1}]$ with $x$, $\int\!\mathrm{d}\mu_{x}$, respectively. We thus observe the advertised reweighting of the MSE term, namely
@@
Next, we perform an EK-type approximation on (69). Namely, we assume $\sigma^{2}$ is large enough compared to $\left(\bm{f}_{1}(x)-\bm{y}_{1}(x)\right)$ to perform a Taylor expansion of the exponent appearing in the action. We note in passing that the next-to-leading order in this expansion Cohen_2021 happens to vanish for $He_{n\geq 5}(x)$ (but not, e.g., for $He_{3}(x)$). Within the leading order expansion, the action in (69) becomes
@@
To test this empirically, we sampled various draws of datasets with $n=100$ from $x\sim{\cal N}[0,1]$, performed GP regression on these with the above kernel using a fixed random test set with $n_{\mathrm{test}}=1000$, and averaged the results over the various draws (see fig. 3). Doing so for $\lambda_{2}=0$ shows an almost zero outcome, as expected. However, taking finite $\lambda_{2}$ reveals a trend that spectral bias Cohen_2021 ; canatar misses. Notably, the fact that our toy example involves only one mode which is integrated out, together with the need to take its variance to be small ($\lambda_{2}/\sigma^{2}\ll 1$) as well as make it unlearnable ($\eta\lambda_{2}/\sigma^{2}\ll 1$) results in a rather small effect. Still, it serves to demonstrate the conceptual phenomena whose importance depends on the setting or the desired accuracy.
Figure 3: Non-Gaussian features and spatial re-weighting effects. Theory versus experiment for the model of sec. V.2 with $n=100$ datapoints $\sigma^{2}=400$ and $\lambda_{>}:=\lambda_{2}=0.1$ (unless stated otherwise). Learning a 5th Hermite polynomial, using a kernel capable of expressing only 1st and 2nd Hermite polynomials should give a zero average predictor (green line) based on the standard theory canatar ; Cohen_2021 . However, due to spatial re-weighting, a coupling between 1st and 5th Hermite polynomial arises leading to a non-zero result. For both $\lambda_{>}=0$ and $\lambda_{>}=0.1$, $m=5$ million trials are performed. The average and standard error (i.e. standard deviation/$\sqrt{m}$) are reported.
VI Discussion
In this work, we applied the powerful formalism of Wilsonian RG to Gaussian Process regression. In particular, treating the output of a trained DNN as a statistical ensemble, we split the associated fluctuations into high- and low-energy modes (based on their stiffness in a random network). Provided the variance of the higher modes is much smaller than the ridge parameter, these may be integrated out in a tractable manner leading to a renormalized GP regression problem. In simple settings where the features associated with those modes are Gaussian, this results in a renormalized ridge parameter. We show that this simple setting can already achieve impressive results on real-world datasets. In particular, we show that our RG formalism can accurately predict the semi power-law scaling behavior of the MSE loss as a function of training dataset size on both MNIST and CIFAR10 regression tasks. In more general settings, one obtains an effective GP regression problem where the ridge parameter becomes spatially dependent. We then applied this to a simple toy model, in which the effects determined via our RG approach are demonstrably missed in the standard eigen-learning framework canatar .